OpenStudy (anonymous):

Solve for c. 3abc + b = 5 c = (5 - b)/(3ab) c = -5/3ab c = 2 - ab\ how does one work this out

OpenStudy (ash2326):

@dirtydan667 \[3abc+b=5\] Subtract b from both the sides, would you try?

OpenStudy (anonymous):

3ac=5....?

OpenStudy (anonymous):

or 3ac=-5b

OpenStudy (ash2326):

ok, I'll explain \[3abc+b=5\] Subtract b from both the sides \[3abc+b-b=5-b\] Now, we see on the left side there is b -b \[3abc+\cancel b-\cancel b=5-b\] \[3abc+0=5-b\] Do you follow this part?

OpenStudy (anonymous):

kinda

OpenStudy (ash2326):

Read it again and let me know if there is a part where you have trouble

OpenStudy (anonymous):

i read it out loud and i get it

OpenStudy (ash2326):

ha ha :D so we have \[3abc=5-b\] Now could you divide both sides by 3ab and post what would you get

OpenStudy (anonymous):

i have no idea where to start with that sir

OpenStudy (ash2326):

Divide both side by 3ab \[\frac{3abc}{3ab}=\frac{5-b}{3ab}\]

OpenStudy (cggurumanjunath):